OpenStudy (anonymous):

Christie has a total of 15 pieces of fruit, all bananas and apples, worth $1.59. Bananas are 13 cents each and apples are 7 cents each. How many bananas and how many apples does she have?

OpenStudy (nurali):

Let 'a' be the amount of apples she has and 'b' the number of bananas. a + b = 15 <----------------- She has 15 pieces of fruit total. .07a + .13b = 1.59 <----- The fruit cost $1.59 total. The first equation can be rewritten b = 15 - a. Now substitute that into the second equation. .07a + .13(15 - a) = 1.59 .07a + 1.95 - .13a = 1.59 -.06a + 1.95 = 1.59 -.06a = -.36 a = -.36 / -.06 a = 6 Now plugging that back into the first equation: b = 15 - 6 = 9.